# On the Design and Analysis of LLM-Based Algorithms

### v2

@@ 3.2 Concrete examples and experiment settings
While the tasks under consideration are motivated by practical scenarios, our study will mostly focus on synthetic task design, like many prior works do. This brings numerous benefits, such as avoiding data contamination, allowing full transparency and control over task configurations, and making the current work as self-contained as possible.
Experiment settings.
We use the following LLMs in our experiments, which cover a wide range of LLM characteristics and inference service:
  • A Llama-3-8B model [Met24], supported by ollama [oll23] and running on a Macbook Pro with a M2 Pro chip and 16GB memory;
  • A Llama-3-70B model [Met24], supported by vLLM [KLZ+23] and running on a server with 4 Nvidia A100-80G GPUs;
  • A GPT-4-Turbo model [Ope24a], accessed via API queries.
All of these LLMs are chat models. Each LLM call involved in our algorithms is prompted in a chat format, based on the sub-task that it is responsible for. Interested readers are referred to the source code for the prompts used in our experiments. We use greedy decoding, which is deterministic, in all experiments.
Below are a few more details about our experiments. (1) For ideal parallelism of LLM calls, we consider parallelism degree $p=4$ and $p=\infty$. Latencies in the presence of parallelism are simulated according to Eq. (3). (2) In all experiments, the number of tokens for a piece of text is estimated using the same tokenizer, namely the cl100k_base encoding of the tiktoken package [^2] https://github.com/openai/openai-cookbook/blob/main/examples/How_to_count_tokens_with_tiktoken.ipynb. This simplification has no effect on the major conclusions from our experiment results. (3) Our experiment results include curves of some error metric (in blue) or cost metric (in red) versus the problem size $n$ or sub-task size $m$. For each curve, we plot the mean and standard deviation of measured metrics from multiple independent trials, i.e. multiple randomly generated task instances.
3.3 Example: counting
As a warm-up exercise, we consider a simple counting task formulated as follows: given a string of length $n$ consisting of letters and digits, the task is to count the number of digits in it. This can be seen as an abstraction or synthetic version of more generic counting tasks in practice.
3.3.1 Algorithm
Below is an LLM-based algorithm that follows the pattern of parallel decomposition:
@@ 3.3.1 Algorithm
    Divide the input string into $k$ disjoint sub-strings of lengths $m_{1},m_{2},\dots,m_{k}$;
@@
    For each $i\in[k]$, let LLM count the number of digits in the $i$-th sub-string, and return its answer $y_{i}$;
  3. 3.
    The final solution of the algorithm is $y=\sum_{i\in[k]}y_{i}$.
@@
For each LLM call, we prompt the LLM to generate its answer without intermediate reasoning steps [^3] On the other hand, step-by-step counting can yield a more accurate answer, but also a significantly higher cost, with $\mathcal{L}_{\textsf{dec}}=O(m)$ rather than $O(1)$.. Thus it is reasonable to assume, based on the LLM’s capability of instruction following, that the text generated by each LLM call has length $\mathcal{L}_{\textsf{dec}}=O(1)$ and can be parsed into a count number.
3.3.2 Analysis
Let us assume for notational convenience that $m_{i}=m$ for all $i\in[k]$, and $k=n/m$ is an integer.
Error metrics.
Denote the ground-truth count for the complete string as $y^{\star}$, and the ground-truth count for the $i$-th sub-string as $y_{i}^{\star}$.
  • Let $\mathcal{E}$ represent the absolute counting error, then
    In other words, the overall error is upper bounded by the sum of errors of sub-tasks.
  • If we let $\mathcal{E}$ represent the normalized counting error instead, then
    $\mathcal{E}(y)\coloneqq\frac{|y-y^{\star}|}{n}\leq\frac{1}{n}\sum_{i\in[k]}|y_{i}-y_{i}^{\star}|=\frac{1}{k}\sum_{i\in[k]}\frac{|y_{i}-y_{i}^{\star}|}{m}=\frac{1}{k}\sum_{i\in[k]}\mathcal{E}(y_{i}).$
    With this metric, the overall error is upper bounded by the average of errors of sub-tasks.
Cost metrics.

### v3

@@ 3.2.3 Retrieval
We present an informal probabilistic analysis for the case of using a Type-B LLM. Given the sub-task size $m$, denote the probability of the first and second failure modes as $p_{1}(m)$ and $p_{2}(m)$ respectively. Then, the success rate of retrieval for the chunk containing the needle is $1-p_{1}(m)$, while the expected number of false positives from the remaining chunks is approximately $k\times p_{2}(m)\asymp n\times p_{2}(m)/m$. One might opt for a relatively small value of $m$, which hopefully increases $1-p_{1}(m)$ and hence mitigates the first failure mode. However, even if $p_{2}(m)/m$ is very small, say $10^{-3}$, the number of false positives can be large if the size $n$ of the original problem is large, causing errors in the solution returned by majority voting.
3.2.4 Summary
With the above analysis, we summarize our answer to Q2 as follows:
Insight 2.
If (1) each sub-task error is monotonely increasing in the sub-task size $m$, and (2) the overall error is bounded by a smooth aggregation (e.g., mean or max) of sub-task errors, then the overall error can be reduced by finer-grained parallel task decomposition, i.e., smaller $m$. Otherwise, caution should be taken: there exist cases (e.g., retrieval with a Type-B LLM) where the final output of the overall algorithm can be sensitive to individual sub-task errors, and finer-grained parallel decomposition (accompanied by a larger number of sub-tasks) does not necessarily imply a smaller final error.
(a) Counting with Llama-3-8B-Instruct; $n=200$ is fixed while $m$ varies.
(b) Sorting with Llama-3-70B-Instruct; $n=200$ is fixed while $m$ varies. See Appendix A.3.3 for analysis and explanation of the zigzag part in the latency with $p=4$.
(c) Retrieval with Llama-3-70B-Instruct; $n=10000$ is fixed while $m$ varies.
(d) Retrieval with Llama-3-8B-Instruct; $n=10000$ is fixed while $m$ varies.
Figure 4: Empirical results for concrete examples of parallel decomposition. Each curve represents the mean and standard deviation of 10 independent trials. See Appendix A for further details and complete results for the case studies on parallel decomposition.
3.3 Empirical validation
The previous analysis has been validated empirically via experiments with Llama 3 models (Meta, 2024). A subset of empirical results are presented in Figure 4, with full implementation details and experimental results deferred to Appendix A. For each figure, we fix the problem size $n$ and use the sub-task size $m$ as the X-axis.
Figure 4 confirms that for all tasks, the end-to-end latency of the overall algorithm is monotonely decreasing in $m$ when the parallelism degree is $p=1$ (i.e., no parallelism), non-monotone in $m$ with a finite parallelism degree $p=4$, and monotonely increasing in $m$ when $p=\infty$, as predicted by Insight 1.
Figure 4 also shows that error metrics increase with $m$ for the counting and sorting tasks, as well as for retrieval with Llama-3-70B-Instruct; in contrast, errors are non-monotone in $m$ for the retrieval task with Llama-3-8B-Instruct, which is susceptible to false positives in retrieval (see Figure 13(b) in Appendix A.4.3) and thus regarded as a Type-B LLM (cf. Section 3.2.3).
4 Case studies and insights: more general patterns
In this section, we demonstrate the proposed analytical framework in action for more diverse patterns of LLM-based algorithms. Our exploration covers unified error analysis under the assumption of additive errors and bounded sensitivity (Section 4.1), two algorithms for iterative retrieval and reasoning that exhibit the pattern of hierarchical decomposition (Section 4.2), and an algorithm that conducts recursive task decomposition and constructs its computational graph dynamically at runtime (Section 4.3). The corresponding analysis and results are summarized in Insights 3, 4 and 5 respectively.
4.1 Unified error analysis for directed acyclic graphs
For generic LLM-based algorithms that are represented by a direct acyclic graphs (DAG) and satisfy the technical assumption below, our framework allows deriving error bounds for them in a unified manner:
Assumption 1.
We say that an LLM-based algorithm satisfies the assumption of additive errors and bounded sensitivity, if for each node $v$ with $k$ inputs $x_{1},\dots,x_{k}$ and a single output $y$, it holds that
for some node-specific additive error $\mathcal{E}_{v}$ and finite sensitivity parameter $S\geq 0$.
Proposition 2.
Consider an LLM-based algorithm, represented by a DAG, that satisfies Assumption 1. Then the error of the output $y(v)$ of any node $v$, including the one that generates the final output of the overall algorithm, is bounded by
$\mathcal{E}\big(y(v)\big)\leq\sum_{w\in\mathcal{V}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow v)}S^{|\text{path}|}\times\mathcal{E}_{w}.$ (6)
Here, $\mathcal{V}$ denotes the node set of the DAG, $|\text{path}|$ denotes the length of a path on the DAG, and $\mathcal{P}(w\rightarrow v)$ represents the set of paths from node $w$ to node $v$ if $w\neq v$, or a singleton set containing one hypothetical path of length 0 if $w=v$. For the special case of $S=1$, this upper bound can be simplified as $\mathcal{E}(y(v))\leq\sum_{w\in\mathcal{V}}|\mathcal{P}(w\rightarrow v)|\times\mathcal{E}_{w}$, where $|\mathcal{P}(w\rightarrow v)|$ denotes the number of paths from $w$ to $v$.
This result, whose proof can be found in Appendix B, characterizes precisely how individual sub-task errors propagate through the graph to the final output of the overall algorithm, depending on the number and lengths of paths between graph nodes and the value of the sensitivity parameter $S$:
Insight 3.
Under the assumption of additive errors and bounded sensitivity, Proposition 2 indicates that: (1) the final error of the overall algorithm is a weighted average of all node-specific additive errors $\{\mathcal{E}_{w}\}$; (2) the impact of each node-specific error $\mathcal{E}_{w}$ grows (resp. shrinks) exponentially along each path to the final output if $S>1$ (resp. $S<1$), and also grows proportionally with the total number of paths.
Justifications for Assumption 1.
In some cases, Assumption 1 can be rigorously proved for a non-LLM node. For example, considering the aggregation node in the counting algorithm, our analysis in Section 3.2.1 implies that Assumption 1 holds true with node-specific additive error $\mathcal{E}_{v}=0$, and sensitivity parameter $S=1$ for the absolute counting error, or $S=1/k$ for the normalized counting error. As another example, considering the aggregation node in the sorting algorithm, Proposition 1 indicates that its output $\ell_{\infty}$ error is bounded by $\mathcal{E}\leq\max\{\mathcal{E}_{1},\dots,\mathcal{E}_{k}\}\leq\mathcal{E}_{1}+\dots+\mathcal{E}_{k}$, which means Assumption 1 holds with $\mathcal{E}_{v}=0$ and $S=1$.
For an LLM node, “additive error and bounded sensitivity” is an assumption that we impose on the stability of an LLM’s behavior when solving a sufficiently simple (with respect to its capability) sub-task. One motivating example comes from our case study to be presented in Section 4.3, which involves many LLM calls that are asked to calculate a 1-Lipschitz function (such as sum or max) of multiple input variables. We have observed empirically that the LLM, with CoT prompting, behaves with sufficient stability and makes small errors only occasionally. It is this property that ensures the final error of the overall algorithm stays low and grows slowly (instead of blowing up) with the task complexity and size of computational graph.
Needless to say, there exist practical situations where Assumption 1 does not hold (e.g., when the LLM’s behavior is sensitive to input errors), in which case Proposition 2 is not applicable.
4.2 Iterative retrieval and reasoning
We consider a task, as an extension of the retrieval task in Section 3.2.3, that requires multi-hop reasoning over multiple clues (“needles”) embedded separately in a long piece of text (“haystack”). For example, the targeted question might be “what is the numeric value of A”, while the needles are “A = B”, “B = C”, and “C = 100”. A practical approach for such a task is iterative retrieval and reasoning (Creswell et al., 2023; Xiong et al., 2024; Qwen-Team, 2024; Yue et al., 2024): at each iteration, the algorithm performs retrieval based on the targeted question and the collected information from previous iterations, followed by reasoning over the retrieved information and deciding whether the targeted question can be answered already or requires further retrieval and reasoning. Within each iteration, retrieval can be done along with chunking, which leads to a pattern of hierarchical decomposition. Two variants of this approach are visualized in Figure 5: the “cyclic” version conducts retrieval over multiple chunks sequentially within each iteration, while the “parallel” version conducts retrieval in parallel. When adopting such an algorithm, two critical design choices need to be made:
@@ 4.2 Iterative retrieval and reasoning
    For the “reasoning” nodes, one might prompt the LLM to “answer directly” for efficiency, or “think step by step” to potentially improve accuracy. Which option is better? How much cost overhead will be caused by chain-of-thought prompting in exchange for accuracy improvement?
@@
    For the task of retrieval from multiple chunks at each iteration, which option would be more efficient, the “cyclic” or “parallel” one?
Our answers, derived with the proposed analytical framework, can be summarized as follows:
Insight 4.
(1) Using chain-of-thought prompting for the reasoning nodes (which play a critical role in deciding the final output) will significantly boost accuracy, while incuring only a minor, additive cost overhead to the overall algorithm. See Eq. (9) and Eq. (10) in Appendix C.3.2 for the quantitative version of this statement. (2) The “cyclic” version requires fewer iterations, while the “parallel” version can benefit from parallelism. As a result, in terms of the end-to-end latency, the “cyclic” version is likely to have advantage when the parallelism degree is small, while the reverse is true otherwise.
These insights have been validated empirically via experiments with Qwen-2-72B-Instruct (Yang et al., 2024). For example, Figure 6(a) shows that, across a broad range of degree $g$ (a measure of task complexity), prompting the LLM to calculate step by step for the reasoning nodes significantly reduces the error of the final output, while incuring negligible or no overhead in the total number of prefilling tokens, and minor overhead in the total number of decoding tokens. Moreover, Figure 6(b) compares the end-to-end latencies of two options for retrieval: “cyclic” achieves lower latencies across a broad range of width $w$ (another measure of task complexity) when there is no parallelism among LLM calls (i.e., $p=1$), wheras “parallel” can benefit from an increased value of parallelism degree $p$ and gain advantages over “cyclic”.
We refer interested readers to Appendix C for full details and results of this case study.
(a) The “cyclic” version.
(b) The “parallel” version.
Figure 5: Two algorithms for iterative retrieval and reasoning. The number of iterations is assumed to be 2 here; in reality, this value is determined adaptively by the algorithm itself at runtime. For clarity, some LLM or non-LLM nodes (cf. Figure 2) are merged into one, and an arrow from the “chunking” node to a shaded block means that each “retrieval” node within the block takes the corresponding chunk as input.
(a) Comparing two prompting methods (“answer directly” versus “step by step”) for the reasoning nodes.
(b) Comparing two versions (“cyclic” versus “parallel”) of iterative retrieval and reasoning.
Figure 6: Empirical results for iterative retrieval and reasoning. The X-axis of each figure denotes a measure of task complexity. See Appendix C for further details and complete results of this case study.
Figure 7: The computational graph of a recursive LLM-based algorithm is constructed dynamically, while a dictionary of solved tasks is maintained to avoid solving the same intermediate sub-task repetitively.
Figure 8: Empirical results for recursive task decomposition. The X-axis of each figure denotes a measure of task complexity. See Appendix D for further details and complete results of this case study.
4.3 Recursive task decomposition
For the last case study, we study LLM-based algorithms with recursive task decomposition. A recursive LLM-based algorithm starts from the original task of concern and recursively generates more intermediate sub-tasks, so that each sub-task can be solved by aggregating the solutions to its children tasks, while being unaware of other sub-tasks involved in the overall algorithm. In particular, solving and/or decomposing each sub-task can be achieved by LLM calls, while the outline of recursive task decomposition remains symbolic. Another feature of a recursive LLM-based algorithm is that its computational graph is not pre-specified or static, but rather constructed dynamically at runtime. Such algorithms have been widely adopted in prior works (Kazemi et al., 2023; Schlag et al., 2023; Prasad et al., 2024; Schroeder et al., 2024; Lee & Kim, 2023; Khot et al., 2023; Zhu et al., 2024).
For concreteness, consider a task that requires finding the value of a specific variable connected to a large number of other variables. Figure 7 shows an example, where the targeted variable A depends on B and C, and variable B further depends on C, D and E, and so on. The LLM-based algorithm need to decide adaptively which variables to query information about, and to execute the necessary calculation along the way. A complete description of the recursive algorithm for this task can be found in Appendix D.2. Despite how different such an algorithm might seem from the previously investigated ones, it can still be covered by our proposed analytical framework:
Insight 5.
(1) The framework proposed in Section 2 is applicable to a recursive LLM-based algorithm, by unrolling it into a directed acyclic graph, as illustrated in Figure 7. (2) For the concrete example, prompting the LLM to conduct arithmetic calculation step by step will significantly boost the accuracy of the overall algorithm, but also increase the cost by a multiplicative factor. See Eq. (11) in Appendix D for the quantitative version of this statement.
The second insight has been validated empirically via experiments with Qwen-2-72B-Instruct (Yang et al., 2024). For example, Figure 8 confirms the accuracy gain by prompting the LLM to calculate step by step, and shows that cost metrics grow linearly with the depth $d$ (a measure of task complexity); furthermore, step-by-step prompting has negligible impacts on the total number of prefilling tokens, but increases the total number of decoding tokens and end-to-end latency multiplicatively.
We refer interested readers to Appendix D for full details and results of this case study.
5 Related works
LLM-based algorithms.
The concept of LLM-based algorithms is fairly broad and general, ranging from a combination of one or multiple LLM calls with some prompt engineering (Yao et al., 2023a; Besta et al., 2023; 2024; Lei et al., 2023; Saha et al., 2023; Zhou et al., 2023; Wang et al., 2023; Prasad et al., 2024; Chang et al., 2024), to LLM-powered agent systems (Mialon et al., 2023; Wu et al., 2023; Hong et al., 2024; Gao et al., 2024a; Shen et al., 2023; Zhuge et al., 2024; Pezeshkpour et al., 2024; Xi et al., 2023; Qian et al., 2024; Kapoor et al., 2024; Sumers et al., 2024; Anthropic, 2024) and compound AI systems (Zaharia et al., 2024; Chen et al., 2024) that augment LLMs with additional abilities like tool use and long-term memory, and to the emerging paradigm of LLM programming (Schlag et al., 2023; Khot et al., 2023; Khattab et al., 2024; Zheng et al., 2024; Kambhampati et al., 2024). Some elements of the framework proposed in Section 2, such as the task decomposition principle, the computational graph representation, evaluation of and comparison between LLM-based algorithms with accuracy and efficiency taken into account simultaneously (Wang et al., 2024; Kapoor et al., 2024), etc., have already appeared in one way or another in these prior works. It is primarily our unified, systematic and formal investigation into the design of generic LLM-based algorithms, and the analysis of their accuracy and efficiency (akin to analysis of traditional algorithms), that distinguish the current work from this vast literature. It is worth noting that (Wu et al., 2025) presents formal analysis for chain-of-thought reasoning, and (Snell et al., 2024) presents formal analysis for sequential and parallel sampling. Both works are similar in spirit to the current work, although the problem formulations and methodologies are mostly orthogonal and complementary.
LLM test-time scaling.
Recent works have started to investigate, analytically or empirically, the scaling laws of LLM test-time computation. One line of research is about the scaling properties of repeated sampling, e.g., randomly sampling multiple generations for the same prompt and then aggregating them for the final output (Chen et al., 2024; Brown et al., 2024; Snell et al., 2024; Wu et al., 2024; Chen et al., 2025), which might be regarded as a stochastic variant of parallel decomposition investigated in Section 3. Our work is orthogonal and complementary to this line, as our analysis is deterministic and targeted at generic patterns of LLM-based algorithms. One potential direction for expanding the analytical framework proposed in this work would be to augment it with stochastic decoding and repeated sampling, along with relevant theoretical results. Another line of works is about improving the output quality of an LLM call by generating more tokens autoregressively, e.g., via chain-of-thought (Wei et al., 2022b; Kojima et al., 2022) reasoning. This idea has been investigated theoretically (Feng et al., 2023; Merrill & Sabharwal, 2024; Li et al., 2024b), and further popularized recently by OpenAI-o1 (OpenAI, 2024a), DeepSeek-R1 (DeepSeek-AI et al., 2025), among many others. As this approach of scaling up the test-time computation of individual LLM calls is becoming more widely adopted, it is important to understand, analytically and quantitatively, how it will impact the overall accuracy and efficiency when such LLM calls are embedded within an LLM-based algorithm. Our work can be useful in achieving this, as illustrated in the case studies in Sections 4.2 and 4.3, where the impacts of prompting LLM calls to “think step by step” versus “answer directly” can be characterized analytically with our proposed framework.
Connections to non-LLM literature.
Our work has drawn inspiration from the area of learning-augmented algorithms (Mitzenmacher & Vassilvitskii, 2022; Lindermayr & Megow, 2022). One standard paradigm of this area is to consider a specific task (e.g., sorting (Bai & Coester, 2023) or clustering (Ergun et al., 2022)), assume access to a black-box machine learning model that satisfies certain properties (e.g., what additional computation or information it can offer), propose a novel algorithm that leverages this ML model, provide theoretical guarantees for its accuracy and efficiency, and show the improvements over traditional, purely symbolic algorithms. Our work is similar in spirit to this line of research, but also substantially different, in that our framework is targeted at generic tasks and LLM-based algorithms, with assumptions on the capabilities, limitations and inference service of LLMs (regarded as general-purpose problem solvers) that are quite different from typical assumptions for task-specific ML models in the literature of learning-augmented algorithms.
Our work is also connected to the literature of distributed computing (Lynch, 1996) (in particular, the MapReduce model (Dean & Ghemawat, 2008; Karloff et al., 2010) that the parallel decomposition pattern in Section 3 closely resembles), as well as classical / non-LLM algorithm design and analysis (Cormen et al., 2022; Roughgarden, 2021). When positioned in this context, one major contribution of our work is to (1) identify the relevance of foundational elements in computer science — task decomposition, computational graphs, MapReduce, etc. — to the LLM context, and (2) integrate LLM-based algorithms with these elements, other theoretical tools, and symbolic programs, via appropriate abstractions and characterization for black-box LLMs. This is the key to facilitating formal analysis for the accuracy and efficiency of LLM-based algorithms. The unique and diverse features of LLMs — and their implications — distinguish our work from the classical / non-LLM literature.
6 Conclusion
This work has introduced an analytical framework for studying the design and analysis of LLM-based algorithms. With the task decomposition principle, the computational graph representation, and other key abstractions tailored to LLMs, we are able to provide formal analysis for the accuracy and efficiency of generic LLM-based algorithms, as demonstrated through a diverse range of case studies.
One limitation of this work is the lack of empirical studies in more realistic settings. Although we have presented case studies and empirical validation for various algorithms and synthetic settings (ranging from the most basic setting — parallel decomposition — to more complex ones that resemble the settings in prior empirical works), there is obviously a gap between them and the most complex agentic workflows in real-world scenarios. Nonetheless, we believe that the results in this work can have broader implications in more general scenarios than the synthetic ones. For example, Insights 1 and 2 can, at the minimum, warn practitioners against the common misconception that “finer-grained task decomposition implies higher accuracy / higher total cost for the overall algorithm”, beyond the parallel decomposition setting where these insights were originally derived. Moreover, we anticipate that our analysis for simple patterns could be used as building blocks for analysis of more complex algorithms since, as highlighted in Anthropic (2024), the most successful LLM agents in industry are often built with “simple, composable patterns”.
We hope that our work will inspire future work on enhancing the analytical framework (e.g., by incorporating certain properties and challenges of realistic LLM usage that we might have abstracted away in this work), or leveraging it to derive more novel and generalizable insights that can assist the best practice of real-world LLM applications.
Acknowledgments
The authors would like to thank the reviewers and Action Editor for their constructive feedback that helps improve this work. We also thank Yuchang Sun for helpful discussion, and the AgentScope team (especially Xuchen Pan, Dawei Gao and Weirui Kuang) for infrastructure support.
Appendix A Supplementary materials for Section 3
This appendix complements Section 3, containing details of experiment settings, additional results for the concrete examples presented in Section 3, and additional examples of parallel decomposition.
A.1 Experiment settings
We use the following LLMs, which cover a wide range of LLM characteristics and inference service:
  • Llama-3-8B-Instruct (Meta, 2024), supported by ollama (ollama, 2023) and running on a Macbook Pro with a M2 Pro chip and 16GB memory;
  • Llama-3-70B-Instruct (Meta, 2024) and Qwen-2-72B-Instruct (Yang et al., 2024), supported by vLLM (Kwon et al., 2023) and running on a server with 4 Nvidia A100-80G GPUs;
  • GPT-4-Turbo (OpenAI, 2024), accessed via API queries.
All of these LLMs are chat models. Each LLM call involved in our algorithms is prompted in a chat format, according to the sub-task that it is responsible for. We use greedy decoding in all experiments.
Below are a few more details about our experiments. (1) For ideal parallelism of LLM calls, we consider parallelism degree $p=4$ and $p=\infty$. Latencies in the presence of parallelism are simulated according to Eq. (3). (2) In all experiments, the number of tokens for a piece of text is estimated using the same tokenizer, namely the cl100k_base encoding of the tiktoken package [^3] https://github.com/openai/openai-cookbook/blob/main/examples/How_to_count_tokens_with_tiktoken.ipynb. This simplification has no effect on the major conclusions from our experiment results. (3) Our experiment results include curves of some error metric (in blue) or cost metric (in red) versus the problem size $n$ or sub-task size $m$. For each curve, we plot the mean and standard deviation of measured metrics from multiple independent trials, i.e., multiple randomly generated task instances.
A.2 Example: counting
A.2.1 Algorithm details
@@
For each LLM call within the counting algorithm presented in Table 1, we prompt the LLM to generate its answer directly without intermediate reasoning steps [^4] On the other hand, step-by-step counting can yield a more accurate answer, but also a significantly higher cost, with $\mathcal{L}_{\textsf{dec}}=O(m)$ rather than $O(1)$.. Thus it is reasonable to assume, given the LLM’s capability of instruction following, that the text generated by each LLM call has length $\mathcal{L}_{\textsf{dec}}=O(1)$ and can be parsed into a count number.
A.2.2 Cost analysis